OpenStudy (anonymous):

A very old vending machine accepts only nickels (n) and dimes (d). Candy costs up to $0.50, but sometimes the machine will dispense candy without any coins being inserted into the machine. Which inequality shows all of the ways to obtain a candy bar from the machine?

OpenStudy (texaschic101):

um...are there answer choices

OpenStudy (anonymous):

ya one sec

OpenStudy (anonymous):

i have no clue what the answer is....

OpenStudy (texaschic101):

I am sorry....I am lost on this one too :(

OpenStudy (anonymous):

ok thnx anyway

OpenStudy (texaschic101):

so we will get some help.. @zzr0ck3r

OpenStudy (texaschic101):

@joemath314159

OpenStudy (texaschic101):

@satellite73

OpenStudy (texaschic101):

@oldrin.bataku

OpenStudy (anonymous):

A very old vending machine accepts only nickels (n) and dimes (d). Candy costs up to $0.50, but sometimes the machine will dispense candy without any coins being inserted into the machine. Which inequality shows all of the ways to obtain a candy bar from the machine? well normally we'd expect our total sum to hit \(50\)cents i.e. \(5n+10d=50\) (which is equivalent to \(n/20+d/10=1/2\) by dividing by \(100\) to 'convert' everything from cents to fractions of dollars). if the machine sometimes dispenses candy without any coins, however, we also expect sometimes \(n=d=0\) to yield a solution, despite \(5(0)+10(0)=0<50\). my guess would then be \(5n+10d\le50\) since it includes all solutions (but more than just that). this is equivalent to \(n/20+d/10\le1/2\)

OpenStudy (anonymous):

oh oops, candy costs UP TO \(50\) cents... so it may be \(\le1/2\) from the get-go

OpenStudy (texaschic101):

wow....I knew some smart person would come...thanks oldrin_bataku

OpenStudy (anonymous):

thnx

OpenStudy (texaschic101):

well..even if I didn't really help with the problem...I tagged and he found your answer. well done :)

OpenStudy (anonymous):

theat's the answer